# Restriction of $p$-modular representations of $U(2,1)$ to a Borel subgroup

### v4

@@ 1 Introduction
$\textnormal{Hom}_{B}(\textnormal{ind}^{G}_{B}\varepsilon,\pi)\cong\textnormal{Hom}_{G}(\textnormal{ind}^{G}_{B}\varepsilon,\pi)$.
@@ 2.2 Weights
Let $\sigma$ be an irreducible smooth representation of $K$. As $K_{1}$ is pro-$p$ and normal in $K$, $\sigma$ factors through the finite group $\Gamma_{K}$, i.e., $\sigma$ is the inflation of an irreducible representation of $\Gamma_{K}$. Conversely, any irreducible representation of $\Gamma_{K}$ inflates to an irreducible smooth representation of $K$. We may therefore identify irreducible smooth representations of $K$ with irreducible representations of $\Gamma_{K}$, and we shall call them weights of $K$ or $\Gamma_{K}$ from now on.
For a weight $\sigma$ of $K$, it is known that $\sigma^{I_{1,K}}$ and $\sigma_{I^{\prime}_{1,K}}$ are one-dimensional, and that the natural composition map $\sigma^{I_{1,K}}\hookrightarrow\sigma\twoheadrightarrow\sigma_{I^{\prime}_{1,K}}$ is an isomorphism of vector spaces ([CE04, Theorem 6.12]). This implies there exists a unique $\lambda_{\beta_{K},\sigma}\in\overline{\mathbf{F}}_{p}$, such that $\beta_{K}\cdot v-\lambda_{\beta_{K},\sigma}v\in\sigma(I^{\prime}_{1,K})$, for $v\in\sigma^{I_{1,K}}$. By [HV15, Proposition 3.16] and the fact that $\beta_{K}\notin I_{K}\cdot I^{\prime}_{K}$, the scalar $\lambda_{\beta_{K},\sigma}$ is zero if $\text{dim}\sigma>1$, and is equal to $\sigma(\beta_{K})$ if $\text{dim}\sigma=1$.
@@ 3 The $I_{1,K}$-invariant maps $S_{K}$ and $S_{-}$
$(1)$.  Let $h\in H_{1}=I_{1,K}\cap H$. Then $S_{K}(hv)=h^{s}\cdot S_{K}v$, for $v\in\pi^{N^{\prime}_{m_{K}}}$, and $S_{-}(hv)=h^{s}\cdot S_{-}v$, , for $v\in\pi^{N_{n_{K}}}$, where $h^{s}$ is short for $\beta_{K}h\beta_{K}$.
@@
This is firstly proved in [Xu17]. We reproduce the arguemnt in full here.
For $(1)$, we note that the group $H_{1}$ acts on $\pi^{N_{n_{K}}}$ and $\pi^{N^{\prime}_{m_{K}}}$, as it normalizes $N_{n_{K}}$ and $N^{\prime}_{m_{K}}$. The statement then follows from the definitions.
@@
A slight variant of $(2)$ holds by the same argument. When $u$ goes through $N_{n_{K}+m}\setminus N_{n_{K}+n}/N_{n_{K}+l}$, the element $u_{1}$ also goes through $N_{n_{K}+m}\setminus N_{n_{K}+n}/N_{n_{K}+l}$, for any $l>n\geq m\geq 0$.
We apply the operators $S_{-}$ and $S_{K}$ to the $I_{1,K}$-invariants of a principal series $\text{ind}^{G}_{B}\varepsilon$. Recall the space $(\text{ind}^{G}_{B}\varepsilon)^{I_{1,K}}$ is spanned by the functions $g_{1}$ and $g_{2}$ which are characterized by: $g_{1}(Id)=1,g_{1}(\beta_{K})=0,g_{2}(Id)=0,g_{2}(\beta_{K})=1$.
@@
Here, $d_{q^{t_{K}}}=\sum_{(x,t)\in L^{\times}_{q^{t_{K}}}}\varepsilon_{0}((h(t)),d_{q^{4-t_{K}}}=\sum_{(x,t)\in L^{\times}_{q^{4-t_{K}}}}\varepsilon_{0}((h(t))$, $\varepsilon_{0}=\varepsilon\mid_{B\cap K}$.
@@ 4.1 Some recaps
$\Psi:C^{\infty}_{c}(N)\rightarrow\kappa_{\varepsilon},f\mapsto\Psi(f),\Psi(f)(b\beta u)=\varepsilon(b)f(u),\forall b\in B,u\in N.$
One verifies easily that $\Psi$ and $\Phi$ are inverse to each other. This gives $C^{\infty}_{c}(N)$ a structure of $B$-representation.
@@
The restriction map induces an isomorphism between the following spaces:
@@
We are done if $\varepsilon\neq\eta\circ\text{det}$ for any character $\eta$. Assume $\varepsilon=\eta\circ\text{det}$ for some $\eta$. After a twist we may assume $\eta=1$. If $\phi\neq 0$, it induces a non-zero map in $\textnormal{Hom}_{B}(1,\pi)\cong\textnormal{Hom}_{G}(1,\pi)$. This in turn implies the map $\phi\in\textnormal{Hom}_{G}(\textnormal{ind}^{G}_{B}1,\pi)$ realizes the trivial character of $G$ as a quotient of $\textnormal{ind}^{G}_{B}1$ (note that $\phi$ is not injective by our assumption), which is not true.
@@
By [Xu18b, Proposition 4.14, 4.15], we have an isomorphism, unique up to a scalar,
@@
2).  By Theorem 4.2 and irreducibility of $\textnormal{ind}^{G}_{B}\varepsilon$, there is an isomorphism $\text{ind}^{G}_{B}\varepsilon\cong\langle G\cdot\phi(\kappa_{\varepsilon})\rangle$.
With these ingredients available, the proof is then identical to that of [Paš07, Corollary 5.5]. ∎
@@
$\textnormal{Hom}_{B}(St,\pi)\cong\textnormal{Hom}_{G}(\textnormal{ind}^{G}_{B}1,\pi)$
@@ 5.2 A key property
Let $\pi$ be an irreducible smooth representation of $G$, and $\sigma$ be a weight of $K$ contained in $\pi$. By [Xu18a, Theorem 1.1], the representation $\pi$ admits Hecke eigenvalues for the spherical Hecke algebra $\mathcal{H}(K,\sigma)\cong\overline{\mathbf{F}}_{p}[T_{\sigma}]$. This implies that the representation $\pi$ is a quotient of $\text{ind}^{G}_{K}\sigma/(T_{\sigma}-\lambda)$, for some scalar $\lambda$. By [Xu18b, Theorem 1.1], the representation $\pi$ is supersingular if $\lambda=0$.
@@
By [Xu18a, Corollary 4.2], there is a non-constant polynomial $P(X)$ such that $\phi\circ P(T_{\sigma})=0$. Assume $P(X)$ is such a polynomial of minimal degree. Take a root $\lambda$ of $P(X)$, and write $P(X)=(X-\lambda)P_{1}(X)$. Put $\phi^{\prime}=\phi\circ P_{1}(T_{\sigma})$. Note that $\phi^{\prime}$ is still a $G$-map from $\textnormal{ind}^{G}_{K}\sigma$ to $\pi$. By our assumption, the map $\phi^{\prime}$ is non-zero and factors through $\textnormal{ind}^{G}_{K}\sigma/(T_{\sigma}-\lambda)$. As $\pi$ is supersingular, we have $\lambda=0$ ([Xu18b, Theorem 1.1]). We conclude $P(X)=X^{n}$ for some $n\geq 1$. ∎
@@
$v=-\alpha\cdot\sum_{u\in(N_{n_{K}}\setminus N_{n_{K}+2})/N_{n_{K}+2}}u\alpha^{-1}v$
@@
$\beta_{K}v=-\sum_{u\in(N_{n_{K}}\setminus N_{n_{K}+2})/N_{n_{K}+2}}\beta_{K}\alpha u\alpha^{-1}v.$ (6)
Applying (1), we see $\beta_{K}\alpha u\alpha^{-1}\in BN^{\prime}_{m_{K}}$, for any $u\in(N_{n_{K}}\setminus N_{n_{K}+2})/N_{n_{K}+2}$. More precisely, for $u=n(\ast,\varpi^{n_{K}}_{E}t)\in N_{n_{K}}\setminus N_{n_{K}+1}$, we have
@@
for $u=n(\ast,\varpi^{n_{K}+1}_{E}t)\in N_{n_{K}+1}\setminus N_{n_{K}+2}$, we have
@@
That gives $\beta_{K}\alpha u\alpha^{-1}v\in\langle B\cdot v\rangle$ for all $u\in(N_{n_{K}}\setminus N_{n_{K}+2})/N_{n_{K}+2}$. We conclude $\beta_{K}v\in\langle B\cdot v\rangle$ from the above equality (6). For our later purpose, we record (6) in a more explicit form:
@@
where, in the first sum $u=n(\ast,\varpi^{n_{K}+2}_{E}t)$ goes through $(N_{n_{K}+2}\setminus N_{n_{K}+3})/N_{n_{K}+4}$; in the second sum $u=n(\ast,\varpi^{n_{K}+1}_{E}t)$ goes through $(N_{n_{K}+1}\setminus N_{n_{K}+2})/N_{n_{K}+2}$. ∎
We proceed to complete the proof of Proposition 5.5. Choose $0\neq v\in\pi^{I_{1,K}}\cap\langle B\cdot w\rangle$ such that $S_{+}v=0$. The above Lemma says $\beta_{K}v\in\langle B\cdot v\rangle$. As $\pi$ is irreducible, we have $\pi=\langle G\cdot v\rangle$. By the Bruit decomposition $G=BI_{1,K}\cup B\beta_{K}I_{1,K}$, we see
@@
The condition in Proposition 5.5 is merely sufficient: for the Steinberg representation $St$, its restriction to $B$ is irreducible, but $S_{+}$ does not annihilate the line $St^{I_{1,K}}$. More precisely, the representation $St$ is defined as $\textnormal{ind}^{G}_{B}1/(1)$, and its $I_{1,K}$-invariants is of one dimension. In the notation of section 3, the space $St^{I_{1,K}}$ is spanned by the image of $g_{1}$. By Proposition 3.5, we check that
@@
We begin with a non-zero vector $v\in\pi^{I_{1,K}}$, and by replacing $v$ by a proper vector in the space $\langle I_{K}\cdot v\rangle$, we may assume $I_{K}$ acts on $v$ by a character $\chi$. As $\pi\mid_{B}$ is irreducible (Theorem 5.5), the map $\phi$ is injective whence $\phi(v)\neq 0$. As $\pi^{\prime}$ is smooth, $\phi(v)$ is fixed by $N^{\prime}_{m_{K}+2m}$ for some $m\geq 0$.
@@
$w=-\sum_{u\in N_{n_{K}}\setminus N_{n_{K}+1}/N_{n_{K}+2}}\beta_{K}u\alpha^{-1}v$
An application of (1) gives that for $u=n(\ast,\varpi^{n_{K}}_{E}t)\in N_{n_{K}}\setminus N_{n_{K}+1}$
@@
$w=-\sum_{u\in N_{n_{K}}\setminus N_{n_{K}+1}/N_{n_{K}+2}}\chi(h(\bar{t}^{-1}))u\alpha^{-1}\cdot v\in\langle B\cdot v\rangle$,
@@
$\phi(w)=-\sum_{u\in N_{n_{K}}\setminus N_{n_{K}+1}/N_{n_{K}+2}}\chi(h(\bar{t}^{-1}))u\alpha^{-1}\cdot\phi(v)$
@@
$\beta_{K}v=-\sum_{u\in N_{n_{K}+1}\setminus N_{n_{K}+1}/N_{n_{K}+2}}\chi(h(\bar{t}^{-1}))u\alpha^{-1}\cdot v$.
@@
$\phi(\beta_{K}v)=-\sum_{u\in N_{n_{K}+1}\setminus N_{n_{K}+1}/N_{n_{K}+2}}\chi(h(\bar{t}^{-1}))u\alpha^{-1}\cdot\phi(v)$
@@
$v\in\langle u\beta_{K}v|u\in N_{n_{K}}\setminus N_{n_{K}+1}/N_{n_{K}+1}\rangle$.
@@
$v=-\sum_{u\in N_{n_{K}}\setminus N_{n_{K}+1}/N_{n_{K}+1}}\chi(h(\bar{t}^{-1}))u\beta_{K}\cdot v$
@@
$\{u\beta_{K}v\mid u\in N_{n_{K}}\setminus N_{n_{K}+1}/N_{n_{K}+1}\}$
Now take a weight $\sigma$ contained in $\tau$, and the unique line $\sigma^{I_{1,K}}$ is given by a non-zero vector which we take as our $w$. Then $w$ is a linear combination of the elements from the above set. This implies, as before that $\phi(w)$ is fixed by $N^{\prime}_{m_{K}+2m-2}$.
@@
where, in the first sum $u=n(\ast,\varpi^{n_{K}+2}_{E}t)$ goes through $(N_{n_{K}+2}\setminus N_{n_{K}+3})/N_{n_{K}+4}$; in the second sum $u=n(\ast,\varpi^{n_{K}+1}_{E}t)$ goes through $(N_{n_{K}+1}\setminus N_{n_{K}+2})/N_{n_{K}+2}$. As $\phi$ is a $B$-map, we firstly see that
$\phi(\beta_{K}w^{\prime})=-\sum_{u}u\alpha^{-2}h(\bar{t}^{-1})\phi(w^{\prime})-\sum_{u}u\alpha^{-1}h(\bar{t}^{-1})\phi(w^{\prime})$
@@
$\beta_{K}\phi(w^{\prime})=-\sum_{u}u\alpha^{-2}h(\bar{t}^{-1})\phi(w^{\prime})-\sum_{u}u\alpha^{-1}h(\bar{t}^{-1})\phi(w^{\prime})$
@@
Recall that the vector $w^{\prime}$ and its image $\phi(w^{\prime})$ are both $I_{1,K}$-invariant. As $G=BI_{1,K}\cup B\beta_{K}I_{1,K}$ and the non-zero vector $w^{\prime}$ generates the representation $\pi$, we conclude that from the above equality that the $B$-map $\phi$ is indeed a $G$-map. ∎

### v5

@@ 1 Introduction
$\textnormal{Hom}_{G}(\textnormal{ind}^{G}_{B}\varepsilon,\pi)\cong\textnormal{Hom}_{B}(\textnormal{ind}^{G}_{B}\varepsilon,\pi)$.
@@ 2.2 Weights
Let $\sigma$ be an irreducible smooth representation of $K$. As $K^{1}$ is pro-$p$ and normal in $K$, $\sigma$ factors through the finite group $\Gamma_{K}$, i.e., $\sigma$ is the inflation of an irreducible representation of $\Gamma_{K}$. Conversely, any irreducible representation of $\Gamma_{K}$ inflates to an irreducible smooth representation of $K$. We may therefore identify irreducible smooth representations of $K$ with irreducible representations of $\Gamma_{K}$, and we shall call them weights of $K$ or $\Gamma_{K}$ from now on.
For a weight $\sigma$ of $K$, it is known that $\sigma^{I_{1,K}}$ and $\sigma_{I^{\prime}_{1,K}}$ are one-dimensional, and that the natural composition map $\sigma^{I_{1,K}}\hookrightarrow\sigma\twoheadrightarrow\sigma_{I^{\prime}_{1,K}}$ is an isomorphism of vector spaces ([CE04, Theorem 6.12]). This implies there exists a unique $\lambda_{\beta_{K},\sigma}\in\overline{\mathbf{F}}_{p}$, such that $\beta_{K}\cdot v-\lambda_{\beta_{K},\sigma}v\in\sigma(I^{\prime}_{1,K})$, for $v\in\sigma^{I_{1,K}}$. By [HV12, Proposition 3.16] and the fact that $\beta_{K}\notin I_{K}\cdot I^{\prime}_{K}$, the scalar $\lambda_{\beta_{K},\sigma}$ is zero if $\text{dim}\sigma>1$, and is equal to $\sigma(\beta_{K})$ if $\text{dim}\sigma=1$.
@@ 3 The $I_{1,K}$-invariant maps $S_{K}$ and $S_{-}$
$(1)$.  Let $h\in H_{0}=I_{K}\cap H$. Then $S_{K}(hv)=h^{s}\cdot S_{K}v$, for $v\in\pi^{N^{\prime}_{m_{K}}}$, and $S_{-}(hv)=h^{s}\cdot S_{-}v$, for $v\in\pi^{N_{n_{K}}}$, where $h^{s}$ is short for $\beta_{K}h\beta_{K}$.
@@
For $(1)$, we note that the group $H_{0}$ acts on $\pi^{N_{n_{K}}}$ and $\pi^{N^{\prime}_{m_{K}}}$, as it normalizes $N_{n_{K}}$ and $N^{\prime}_{m_{K}}$. The statement then follows from the definitions.
@@
A slight variant of $(2)$ holds by the same argument. When $u$ goes through $(N_{n_{K}+m}\smallsetminus N_{n_{K}+n})/N_{n_{K}+l}$, the element $u_{1}$ also goes through $(N_{n_{K}+m}\smallsetminus N_{n_{K}+n})/N_{n_{K}+l}$, for any $l>n\geq m\geq 0$.
We apply the operators $S_{-}$ and $S_{K}$ to the $I_{1,K}$-invariants of a principal series $\text{ind}^{G}_{B}\varepsilon$. The space $(\text{ind}^{G}_{B}\varepsilon)^{I_{1,K}}$ is two dimensional and a basis of that is given by the functions $g_{1}$ and $g_{2}$: the function $g_{1}$ is supported on $BI_{K}$ and satisfies $g_{1}(Id)=1$; the function $g_{2}$ is supported on $B\beta_{K}I_{K}$ and satisfies that $g_{2}(\beta_{K})=1$.
@@
Here, $d_{q^{t_{K}}}=\sum_{n(x,t)\in(N_{n_{K}}\smallsetminus N_{n_{K}+1})/N_{n_{K}+1}}\varepsilon_{0}((h(t))$,
$d_{q^{4-t_{K}}}=\sum_{n(x,t)\in(N_{n_{K}+1}\smallsetminus N_{n_{K}+2})/N_{n_{K}+2}}\varepsilon_{0}((h(t))$, $\varepsilon_{0}=\varepsilon\mid_{B\cap K}$.
@@ 4.1 Some recaps
$\Psi:C^{\infty}_{c}(N)\rightarrow\kappa_{\varepsilon},f\mapsto\Psi(f),\Psi(f)(b\beta u)=\varepsilon(b)f(u),\Psi(f)(b)=0,\forall b\in B,u\in N.$
For $b=hu_{1}\in B$ where $h\in H,u_{1}\in N$, and $f\in C^{\infty}_{c}(N)$, we put $b\cdot f(u)=\varepsilon(h^{s})f((h^{-1}uh)u_{1})$, where $h^{s}$ denotes $\beta h\beta$. This gives $C^{\infty}_{c}(N)$ a structure of $B$-representation. We check easily that $\Psi$ and $\Phi$ are both $B$-equivariant, and are inverse to each other.
@@
Let $\pi$ be any smooth representation of $G$. The restriction map induces an isomorphism between the following spaces:
@@
We are done if $\varepsilon\neq\eta\circ\text{det}$ for any character $\eta$. Assume $\varepsilon=\eta\circ\text{det}$ for some $\eta$. After a twist we may assume $\eta=1$. If $\phi\neq 0$, it induces a non-zero map in $\textnormal{Hom}_{B}(1,\pi)$, which by the argument of Lemma 4.3 is in $\textnormal{Hom}_{G}(1,\pi)$. This implies the map $\phi\in\textnormal{Hom}_{G}(\textnormal{ind}^{G}_{B}1,\pi)$ realizes the trivial character of $G$ as a quotient of $\textnormal{ind}^{G}_{B}1$, which is not true.
@@
By [Xu18, Proposition 4.15, 4.16], we have an isomorphism, unique up to a scalar,
@@
2).  Applying Theorem 4.2 to $\pi^{\prime}=\langle G\cdot\phi(\kappa_{\varepsilon})\rangle$ and using irreducibility of $\textnormal{ind}^{G}_{B}\varepsilon$, there is an isomorphism $\text{ind}^{G}_{B}\varepsilon\cong\langle G\cdot\phi(\kappa_{\varepsilon})\rangle$.
By 1), the map $\phi$ lies in $\textnormal{Hom}_{B}(\textnormal{ind}^{G}_{B}\varepsilon,\langle G\cdot\phi(\kappa_{\varepsilon})\rangle)$ but by 2) the latter space is isomorphic to $\textnormal{End}_{B}(\textnormal{ind}^{G}_{B}\varepsilon)$. We deduce that $\phi$ is $G$-equivariant by the last assertion of Remark 4.4. ∎
@@
$\textnormal{Hom}_{G}(\textnormal{ind}^{G}_{B}1,\pi)\cong\textnormal{Hom}_{B}(St,\pi)$
@@ 5.2 A key property
Let $\pi$ be an irreducible smooth representation of $G$, and $\sigma$ be a weight of $K$ contained in $\pi$. By [Xu25, Theorem 1.1], the representation $\pi$ admits Hecke eigenvalues for the spherical Hecke algebra $\mathcal{H}(K,\sigma)\cong\overline{\mathbf{F}}_{p}[T_{\sigma}]$. This implies that the representation $\pi$ is a quotient of $\text{ind}^{G}_{K}\sigma/(T_{\sigma}-\lambda)$, for some scalar $\lambda$. By definition 5.1, the representation $\pi$ is supersingular if $\lambda=0$.
@@
By [Xu25, Corollary 4.2], there is a non-constant polynomial $P(X)$ such that $\phi\circ P(T_{\sigma})=0$. Assume $P(X)$ is such a polynomial of minimal degree. Take a root $\lambda$ of $P(X)$, and write $P(X)=(X-\lambda)P_{1}(X)$. Put $\phi^{\prime}=\phi\circ P_{1}(T_{\sigma})$. Note that $\phi^{\prime}$ is still a $G$-map from $\textnormal{ind}^{G}_{K}\sigma$ to $\pi$. By our assumption, the map $\phi^{\prime}$ is non-zero and factors through $\textnormal{ind}^{G}_{K}\sigma/(T_{\sigma}-\lambda)$. As $\pi$ is supersingular, we have $\lambda=0$ ([Xu18, Theorem 1.1]). We conclude $P(X)=X^{n}$ for some $n\geq 1$. ∎
@@
$v=-\alpha\cdot\sum_{u\in(N_{n_{K}}\smallsetminus N_{n_{K}+2})/N_{n_{K}+2}}u\alpha^{-1}v$
@@
$\beta_{K}v=-\sum_{u\in(N_{n_{K}}\smallsetminus N_{n_{K}+2})/N_{n_{K}+2}}\beta_{K}\alpha u\alpha^{-1}v.$ (6)
Applying (1), we see $\beta_{K}\alpha u\alpha^{-1}\in BN^{\prime}_{m_{K}}$, for any $u\in(N_{n_{K}}\smallsetminus N_{n_{K}+2})/N_{n_{K}+2}$. More precisely, for $u=n(\ast,\varpi^{n_{K}}_{E}t)\in N_{n_{K}}\smallsetminus N_{n_{K}+1}$, we have
@@
for $u=n(\ast,\varpi^{n_{K}+1}_{E}t)\in N_{n_{K}+1}\smallsetminus N_{n_{K}+2}$, we have
@@
That gives $\beta_{K}\alpha u\alpha^{-1}v\in\langle B\cdot v\rangle$ for all $u\in(N_{n_{K}}\smallsetminus N_{n_{K}+2})/N_{n_{K}+2}$. We conclude $\beta_{K}v\in\langle B\cdot v\rangle$ from the above equality (6). For our later purpose, we record (6) in a more explicit form:
@@
where, in the first sum $u=n(\ast,\varpi^{n_{K}+2}_{E}t)$ goes through $(N_{n_{K}+2}\smallsetminus N_{n_{K}+3})/N_{n_{K}+4}$; in the second sum $u=n(\ast,\varpi^{n_{K}+1}_{E}t)$ goes through $(N_{n_{K}+1}\smallsetminus N_{n_{K}+2})/N_{n_{K}+2}$. ∎
We proceed to complete the proof of Proposition 5.5. Choose $0\neq v\in\pi^{I_{1,K}}\cap\langle B\cdot w\rangle$ such that $S_{+}v=0$. The above Lemma says $\beta_{K}v\in\langle B\cdot v\rangle$. As $\pi$ is irreducible, we have $\pi=\langle G\cdot v\rangle$. By the Bruhat decomposition $G=BI_{1,K}\cup B\beta_{K}I_{1,K}$, we see
@@
The condition in Proposition 5.5 is merely sufficient: for the Steinberg representation $St$, its restriction to $B$ is irreducible, but $S_{+}$ does not annihilate the line $St^{I_{1,K}}$. More precisely, the representation $St$ is defined as $\textnormal{ind}^{G}_{B}1/(1)$, and its $I_{1,K}$-invariants is of one dimension. In the notation of section 3, the space $St^{I_{1,K}}$ is spanned by the image of the function $g_{1}$. By Proposition 3.5, we check that
@@
We begin with a non-zero vector $v\in\pi^{I_{1,K}}$, and by replacing $v$ by some other vector in the space $\langle I_{K}\cdot v\rangle$, we may assume $I_{K}$ acts on $v$ by a character $\chi$. As $\pi\mid_{B}$ is irreducible (Theorem 5.5), the map $\phi$ is injective whence $\phi(v)\neq 0$. As $\pi^{\prime}$ is smooth, $\phi(v)$ is fixed by $N^{\prime}_{m_{K}+2m}$ for some $m\geq 0$.
@@
$w=-\sum_{u\in(N_{n_{K}}\smallsetminus N_{n_{K}+1})/N_{n_{K}+2}}\beta_{K}u\alpha^{-1}v$
An application of (1) gives that for $u=n(\ast,\varpi^{n_{K}}_{E}t)\in N_{n_{K}}\smallsetminus N_{n_{K}+1}$
@@
$w=-\sum_{u\in(N_{n_{K}}\smallsetminus N_{n_{K}+1})/N_{n_{K}+2}}\chi(h(\bar{t}^{-1}))u\alpha^{-1}\cdot v\in\langle B\cdot v\rangle$,
@@
$\phi(w)=-\sum_{u\in(N_{n_{K}}\smallsetminus N_{n_{K}+1})/N_{n_{K}+2}}\chi(h(\bar{t}^{-1}))u\alpha^{-1}\cdot\phi(v)$.
@@
$\beta_{K}v=-\sum_{u\in(N_{n_{K}+1}\smallsetminus N_{n_{K}+2})/N_{n_{K}+2}}\chi(h(\bar{t}^{-1}))u\alpha^{-1}\cdot v\in\langle B\cdot v\rangle$.
@@
$\phi(\beta_{K}v)=-\sum_{u\in(N_{n_{K}+1}\smallsetminus N_{n_{K}+2})/N_{n_{K}+2}}\chi(h(\bar{t}^{-1}))u\alpha^{-1}\cdot\phi(v)$.
@@
$v\in\langle u\beta_{K}v|u\in(N_{n_{K}}\smallsetminus N_{n_{K}+1})/N_{n_{K}+1}\rangle$.
@@
$v=-\sum_{u\in(N_{n_{K}}\smallsetminus N_{n_{K}+1})/N_{n_{K}+1}}\chi(h(\bar{t}^{-1}))u\beta_{K}\cdot v$
@@
$\{u\beta_{K}v\mid u\in(N_{n_{K}}\smallsetminus N_{n_{K}+1})/N_{n_{K}+1}\}$
Now take a weight $\sigma$ contained in $\tau$, and the unique line $\sigma^{I_{1,K}}$ is given by a non-zero vector which we take as our $w$. Then $w$ is a linear combination of the elements from the above set. This implies, as before that $\phi(w)$ is fixed by $N^{\prime}_{m_{K}+2m-2}$ [^1] Note that in this case $w$ can be chosen as $v$, but as the argument indicates there are other possibilities. This means that, under the assumption $S_{-}v=S_{K}v=0$, the vector $\phi(v)$ is already fixed by the whole group $I_{1,K}$..
@@
where, in the first sum $u=n(\ast,\varpi^{n_{K}+2}_{E}t)$ goes through $(N_{n_{K}+2}\smallsetminus N_{n_{K}+3})/N_{n_{K}+4}$, which we denote by $\mathfrak{N}_{1}$; in the second sum $u=n(\ast,\varpi^{n_{K}+1}_{E}t)$ goes through $(N_{n_{K}+1}\smallsetminus N_{n_{K}+2})/N_{n_{K}+2}$, which we denote by $\mathfrak{N}_{2}$. As $\phi$ is a $B$-map, we firstly see that
$\phi(\beta_{K}w^{\prime})=-\sum_{u\in\mathfrak{N}_{1}}u\alpha^{-2}h(\bar{t}^{-1})\phi(w^{\prime})-\sum_{u\in\mathfrak{N}_{2}}u\alpha^{-1}h(\bar{t}^{-1})\phi(w^{\prime})$
@@
$\beta_{K}\phi(w^{\prime})=-\sum_{u\in\mathfrak{N}_{1}}u\alpha^{-2}h(\bar{t}^{-1})\phi(w^{\prime})-\sum_{u\in\mathfrak{N}_{2}}u\alpha^{-1}h(\bar{t}^{-1})\phi(w^{\prime})$
@@
Recall that the vector $w^{\prime}$ and its image $\phi(w^{\prime})$ are both $I_{1,K}$-invariant. As $G=BI_{1,K}\cup B\beta_{K}I_{1,K}$ and the non-zero vector $w^{\prime}$ generates the representation $\pi$, we conclude from the above equality that the $B$-map $\phi$ is indeed a $G$-map. ∎